OpenStudy (anonymous):

Jack earned twice as much as Mary and Jill earned $8 more than Mary. Who made the most money?   Jack Mary Jill Can't Answer

OpenStudy (anonymous):

Let Jack = A, Mary = M, Jill = J A = 2M, J = 8+M By looking at this relation, Mary made the least, and Jack made the most.

OpenStudy (anonymous):

Can't actually answer that one yet. If Mary earns less than $8 Jill wins - more than $8 Jack wins. Mary earns say $2: Jill gets 2+8=10; Jack gets 2x2=4. If Mary got 10: Jill gets 10+8=18; Jack gets 2x10=20. If Mary earned exactly 8 - Mary and Jack both get $16. So, unless you know Mary's starting pay there is no way to know who got the most.